OpenStudy (anonymous):

Find the value of c when a=1 and b=7

OpenStudy (angela210793):

is there any function given?

OpenStudy (anonymous):

complete the exercise on the equation a=c/2b

OpenStudy (angela210793):

1=c/14-->c=14

OpenStudy (angela210793):

14=2b=2*7

OpenStudy (anonymous):

cross multiply c=2ab c=2*1*7=14

OpenStudy (anonymous):

can you guys explain how you got your answers?

OpenStudy (angela210793):

all u do is replace a with 1 and b with 7 so y'll have 1=c/(2*7)=1/14 multiply both sides with 14 and u get 14=c